# Quantum Doubly Stochastic Transformers

### v1

@@ 6 Conclusion
Here, we proposed the quantum doubly stochastic Transformer. We conceived this method by connecting the centerpiece of a novel variational quantum circuit (devised for optimal transport [37]) with the Transformer, facilitated through the empirical observation by Sander et al. [42] that doubly-stochastic attention improves performance in Transformers. By extending the QontOT circuit from scalars to matrices we enabled its integration into a ViT, thus providing the first parametric, doubly-stochastic Transformer. Notably, the QDSFormer presents a meta-class of Sinkformers because it estimates DSMs parametrically, i.e., it can be optimized to learn arbitrary transformations onto the Birkhoff polytope. Moreover, to our knowledge, there is no classical, parametric approach to estimate DSMs, thus the QDSFormer is a promising candidate for hybrid quantum-classical neural networks trained on quantum hardware.
@@
Our empirical expressivity analysis revealed that the quantum circuit produces DSMs that are more diverse, preserve information better and have higher entropy than DSMs from Sinkhorn’s algorithm. On multiple simple object recognition tasks, the QDSFormer exhibited significantly higher accuracy, outperforming a ViT and a Sinkformer in most cases. Our usage of quantum attention substantially stabilizes the notoriously unstable ViT training on small-scale data, as evidenced by the performance on a compositional object recognition task (Figure 5), previously used to study ViT training dynamics [24]. Albeit these results are promising, all experiments were performed on comparably small-scale, due to the currently poor scaling of quantum computers in general (which is expected to improve). Notably, by leveraging QR decomposition, we also proposed a novel, quantum-inspired attention flavor. Broadly speaking, outsourcing the activation function to a parametric quantum circuit might be seen a computational overhead, however, we envision that this may reveal potential benefits (typically in small-data, small-model and short-training settings [1]) that are out of reach for classical hardware. To that end, future work could explore concurrent optimization of ViT and circuit parameters via the parameter-shift rule on real quantum hardware.
Appendix A Circuit execution times
@@
Figure A1: QontOT circuit execution times for DSM of size $4$ (A) and $8$ (B) for different combinations of qubits and circuit layers. C and D show the relative increase in execution time as a function of increasing the number of layers (C), and qubits (D). Adding more layers has a sublinear effect on runtime, adding qubits requires exponential more runtime. The minimal number of auxiliary qubits is $\log_{2}(n)+1$ and the total number of qubits is $2(q_{d}+q_{a})$ where $q_{d}$ and $q_{a}$ are data and auxilliary qubits respectively.
Appendix B Empirical circuit expressivity
@@
Appendix C QDSFormer results
C.1 Hyperparameters
For experiments on MNIST, FashionMNIST and the seven MedMNIST datasets, the ViT was configured with a hidden dimension of 128 and an MLP dimension expansion factor of 1. The model was tested with 1 to 4 Transformer layers, each containing a single attention head. No dropout was applied, and the batch size was set to 100. For the optimizer, Adam [31] was used, and the learning rate schedule followed the setup in Sander et al. [42], with an initial learning rate of 5e-4, decreasing by a factor of 10 at epochs 31 and 45. For the more complex Eureka dataset, comprised of 56x56 RGB images, the hidden dimension was increased to 256, and a larger batch size of 512 was used. The MLP expansion factor was also doubled to 2. A cosine learning rate schedule was used with the optimizer AdamW [35]. The scheduler uses 5 warmup epochs with a warm-up learning rate of 1e-6, the decay rate is set to 0.1 and the minimum learning rate is 1e-5, the other parameters follows the default TIMM settings [56]. For the optimizer the weight decay is 0.05 and betas $(0.9,0.999)$. For the Eureka dataset, no Exponential Moving Average (EMA) is used.
C.2 QontOT ansatz types
Table A1: Ablation study for a $2$-layer QDSFormer with different circuit ansatz types and varying number of layers on FashionMNIST. Mean/std of $5$ runs.
@@ C.2 QontOT ansatz types
8 $\mathbf{89.9}_{\pm 0.15}$ $89.8_{\pm 0.15}$ $89.4_{\pm 0.17}$ $88.4_{\pm 0.20}$
Simple: This ansatz is the most generic and resembles a checkerboard structure formed by 4-parameter unit-blocks acting on two qubits each [30, 36]. If all parameters are zero, it falls back to the identity. This ansatz is convenient because it is shallow in simulation but whose depth may vary depending on qubit layout of the quantum hardware.
Parted: This ansatz partitions the Simple ansatz into two parts: $U=U_{1}\otimes U_{2}$, where $U_{2}$ operates normally, and $U_{1}$ is transposed and placed around the initial Bell state. This design reduces the original Simple ansatz circuit depth nearly by half, which may be more efficient on certain quantum hardware. However depending on the qubit layout of the quantum hardware, it carries the potential of increased transpiled circuit depth, as the 2-qubit gates may act on distant qubits necessitating additional swap gates upon transpilation, which we observed on IBM Eagle and Heron quantum processing units. Unless mentioned otherwise, we used this ansatz in all our experiments as it yields shallower circuits in simulation and the increased depth compared to the Simple ansatz was negligible at tested system sizes.
Centrosymmetric: This was the predominantly used ansatz by Mariella et al. [37]. It is less generic, biasing toward properties of centrosymmetric matrices.
Trotter: This ansatz implements a second-order Trotter decomposition [47]. Each circuit layer corresponds to a Trotter step.
C.3 Time series classification
Table A2: Micro-F1 on IR spectra dataset across $5$-fold cross-validation with a 1-layer ViT. QDSFormer uses 16 circuit layers for both DSM sizes.
@@ C.5 Differentiable circuit
Appendix D Counting DSMs
In Section D.1 we provide a full analytical solution to count the number of $3\times 3$ DSMs for a given discretization $p\in\mathbb{N}$. In Section D.2 we strive to extend the analytic solution to an arbitrary $n\in\mathbb{N}$ but only provide a partial solution. Finally, in Section D.3 we provide numerical results from brute-force calculation of the number of DSMs that verified the explicit analytical solution in Section D.1.
D.1 Analytical solution for $n=3$
@@ D.2.1 Constraint 2
The sum of the $n-1\times n-1$ inner matrix must not be below $n-2$ [10].
@@
unique possible sums. These sums are regularly spaced from 00 to $(n-1)^{2}$ with a step size of $p-1$, i.e., $u_{i}=\left\{\frac{i}{p-1}\mid i\in\{0,1,...,|u|\}\right\}$. This allows conversion to an integer problem (by multiplication of $p-1$) and apply Stars & Bars Theorem 2.
@@
Specifically, we set $k:=(n-1)^{2}$ and then define the set of sums that violate the constraint as $S:=\{s\in\mathbb{N}_{0}\mid 0\leq s\leq(n-2)(p-1)\}$. Thus $|S|=(n-2)(p-1)$ We then compute the violations via:
@@
where $\binom{(n-1)^{2}}{m}$ accounts for the number of ways to choose $m$ out of $(n-1)^{2}$ variables that exceed $p-1$.
@@
$c2(n,p)=\sum_{s=0}^{(n-2)(p-1)}c_{2}(n,p,s)\\ =\sum_{s=0}^{|S|}\sum_{m=0}^{(n-1)^{2}}(-1)^{m}\binom{(n-1)^{2}}{m}\begin{cases}0&\text{if }s-mp\leq 0,\\ \binom{s-mp+n^{2}-2n}{n^{2}-2n}&\text{else}.\end{cases}$ (21)
where $|S|=(n-2)(p-1)$.

### v2

@@ 6 Conclusion
Here, we proposed the Quantum Doubly Stochastic Transformer. We conceived this method by connecting the centerpiece of a novel variational quantum circuit [mariella2024quantum] with the Transformer, facilitated through the empirical observation that doubly-stochastic attention improves performance in Transformers [sander2022sinkformers]. By extending the QontOT circuit from scalars to matrices we enabled its integration into a ViT, thus providing the first parametric, doubly-stochastic Transformer. Notably, the QDSFormer presents a meta-class of Sinkformers because it estimates DSMs parametrically, i.e., it can be optimized to learn arbitrary transformations onto the Birkhoff polytope. Moreover, to our knowledge, there is no classical, parametric approach to estimate DSMs, thus the QDSFormer is a promising candidate for hybrid quantum-classical neural networks trained on quantum hardware.
@@
Our empirical expressivity analysis revealed that the quantum circuit produces DSMs that are more diverse, preserve information better and have higher entropy than DSMs from Sinkhorn’s algorithm. On multiple simple object recognition tasks, the QDSFormer exhibited significantly higher accuracy, outperforming a ViT and a Sinkformer in most cases. Our usage of quantum attention substantially stabilizes the notoriously unstable ViT training on small-scale data, as evidenced by the performance on a compositional object recognition task (4(b)), previously used to study ViT training dynamics [hoffmann2024eureka]. Albeit these results are promising, all experiments were performed on comparably small-scale, due to the currently poor scaling of quantum computers in general (which is expected to improve). Notably, by leveraging QR decomposition, we also proposed a novel, quantum-inspired attention flavor. Broadly speaking, outsourcing the activation function to a parametric quantum circuit might be seen a computational overhead, however, we envision that this may reveal potential benefits (typically in small-data, small-model and short-training settings [abbas2021power]) that are out of reach for classical hardware. To that end, future work could explore concurrent optimization of ViT and circuit parameters via the parameter-shift rule on real quantum hardware.
Funding Declaration
This project received no external funding.
Competing Interests
The authors declare no competing financial or non-financial interests.
Author Contributions.
Conceptualization: J.B.; Project administration: J.B.; Supervision: J.B.; Validation: J.B.;
J.B.F.SK.RF.U.N.W.A.S.ConceptualizationProject administrationSupervisionValidation
Appendix A Quantum hardware experiment
We measured the extent and the effect of quantum hardware noise on a DSM produced by our quantum circuit. To that end, we picked a shallow circuit ($1$ layer) with $14$ qubits and computed the ground-truth $8\times 8$ DSM for a random input matrix through statevector simulation. We then transpiled the circuit on three different quantum computers (Cusco, Brisbane and Torino, for details see Table A1) available to the public via the IBM Quantum Platform.
Table A1: Comparison of IBM quantum backends (EPLG = Error Per Layer Gate).
System Architecture Qubits EPLG (%)
Torino Heron R1 133 1.3
Brisbane Eagle R3 127 2.2
Cusco Eagle R3 127 6.8
After using transpilation optimization level $1$, we obtained a circuit with a 2-qubit-depth of $15$ and and a total of $52$ two-qubit gates. As error mitigation techniques, we used dynamical decoupling [ezzell2023dynamical], Pauli twirling [wallman2016noise] and a projection to the Birkhoff polytope of the approximate-DSM obtained from the quantum circuit (see Section 2.3.2). The results, shown in Figure A1, indicate that, consistently, Cusco was the noisiest machine and Torino yielded the best results. Moreover, in general, beyond a shot count of $10,000$ little performance improvement can be observed. This is a positive finding because it is substantially below the theoretical minimum given by the shot noise limit ($640,000$). However, the deviation from the exact DSM, measured in Frobenius Distance, was substantial (Figure A1A). We analyzed the root cause of this and found that the deviation can be largely attributed to an increase in entropy. DSMs obtained from noisy quantum hardware converge toward the center of the Birkhoff polytope ($1/n$ in every cell). The relative ordering of the absolute values instead is largely preserved (Figure A1C).
Figure A1: Hardware experiment on different quantum computers available via IBM Quantum Platform. A The Frobenius distance between the hardware-obtained DSM to its noise-free equivalent. B: The spearman rank correlation between the $64$ values in the noise-free and hardware-obtained DSMs show that the ordering of values is largely preserved. Statevectorsampler here denotes finite sampling from an ideal, noise-free statevector.
Appendix B Circuit execution times
@@
Figure A2: QontOT circuit execution times for DSM of size 44 (A) and 88 (B) for different combinations of qubits and circuit layers. C and D show the relative increase in execution time as a function of increasing the number of layers (C), and qubits (D). Adding more layers has a sublinear effect on runtime, adding qubits requires exponential more runtime. The minimal number of auxiliary qubits is $\log_{2}(n)+1$ and the total number of qubits is $2(q_{d}+q_{a})$ where $q_{d}$ and $q_{a}$ are data and auxilliary qubits respectively.
Appendix C Empirical circuit expressivity
@@
Appendix D QDSFormer results
D.1 Hyperparameters
For experiments on MNIST, FashionMNIST and the seven MedMNIST datasets, the ViT was configured with a hidden dimension of 128 and an MLP dimension expansion factor of 1. The model was tested with 1 to 4 Transformer layers, each containing a single attention head. No dropout was applied, and the batch size was set to 100. For the optimizer, Adam [Kingma2014AdamAM] was used, and the learning rate schedule followed the setup in sander2022sinkformers, with an initial learning rate of 5e-4, decreasing by a factor of 10 at epochs 31 and 45. For the more complex Eureka dataset, comprised of 56x56 RGB images, the hidden dimension was increased to 256, and a larger batch size of 512 was used. The MLP expansion factor was also doubled to 2. A cosine learning rate schedule was used with the optimizer AdamW [loshchilov2018adamW]. The scheduler uses 5 warmup epochs with a warm-up learning rate of 1e-6, the decay rate is set to 0.1 and the minimum learning rate is 1e-5, the other parameters follows the default TIMM settings [rw2019timm]. For the optimizer the weight decay is 0.05 and betas $(0.9,0.999)$. All studied imaging datasets (MNIST, FashionMNIST and seven types of MedMNIST datasets) come with predefined train/validation/test splits. On the infrared spectral data of molecules from alberts2025unraveling we performed a 5-fold cross validation with $80\%/20\%$ train/test split. Hyperparameters were kept identical to previous experiments. For the Eureka dataset, no Exponential Moving Average (EMA) is used. Experiments were conducted on POWER8 infrastructure in Python 3.9 with PyTorch [paszke2017automatic] $1.13.1$ on machines with $16$ cores of 32GiB RDIMM DDR4 2.7 GHz. Due to the small size of the ViTs, training took between few hours and a day (for the slowest, i.e., end-to-end-differentiable configuration of the QDSFormer). The Sinkformer [sander2022sinkformers] and the standard ViT implementation are taken from the original author’s repository: https://github.com/michaelsdr/sinkformers. The results on the compositional Eureka dataset [hoffmann2024eureka] were generated with the ViT implementation of the original authors: https://github.com/boschresearch/eurekaMoments. The implementation of the QontOT circuit was implemented as described in mariella2024quantum and, as described in the main text, adapted to digest matrix (or vector) inputs rather than scalars only.
D.2 QontOT ansatz types
Table A2: Ablation study for a 22-layer QDSFormer with different circuit ansatz types and varying number of layers on FashionMNIST. Mean/std of 55 runs.
@@ D.2 QontOT ansatz types
8 $\mathbf{89.9}_{\pm 0.15}$ $89.8_{\pm 0.15}$ $9.4_{\pm 0.17} 88$$8.4_{\pm 0.20}$
Simple: This ansatz is the most generic and resembles a checkerboard structure formed by 4-parameter unit-blocks acting on two qubits each [khatri2019quantum, madden2022best]. If all parameters are zero, it falls back to the identity. This ansatz is convenient because it is shallow in simulation but whose depth may vary depending on qubit layout of the quantum hardware.
Parted: This ansatz partitions the Simple ansatz into two parts: $U=U_{1}\otimes U_{2}$, where $U_{2}$ operates normally, and $U_{1}$ is transposed and placed around the initial Bell state. This design reduces the original Simple ansatz circuit depth nearly by half, which may be more efficient on certain quantum hardware. However depending on the qubit layout of the quantum hardware, it carries the potential of increased transpiled circuit depth, as the two-qubit gates may act on distant qubits necessitating additional swap gates upon transpilation, which we observed on IBM Eagle and Heron quantum processing units. Unless mentioned otherwise, we used this ansatz in all our experiments as it yields shallower circuits in simulation and the increased depth compared to the Simple ansatz was negligible at tested system sizes.
Centrosymmetric: This was the predominantly used ansatz by mariella2024quantum. It is less generic, biasing toward properties of centrosymmetric matrices.
Trotter: This ansatz implements a second-order Trotter decomposition [smith2019simulating]. Each circuit layer corresponds to a Trotter step.
D.3 Time series classification
Table A3: Micro-F1 on IR spectra dataset across 55-fold cross-validation with a 1-layer ViT. QDSFormer uses 16 circuit layers for both DSM sizes.
@@ D.5 Differentiable circuit
Appendix E Counting DSMs
In Section E.1 we provide a full analytical solution to count the number of $3\times 3$ DSMs for a given discretization $p\in\mathbb{N}$. In Section E.2 we strive to extend the analytic solution to an arbitrary $n\in\mathbb{N}$ but only provide a partial solution. Finally, in Section D.3 we provide numerical results from brute-force calculation of the number of DSMs that verified the explicit analytical solution in Section E.1.
E.1 Analytical solution for $n=3$
@@ E.2.1 Constraint 2
The sum of the $n-1\times n-1$ inner matrix must not be below n2n-2 [chan1999volume].
@@
unique possible sums. These sums are regularly spaced from 0 to $(n-1)^{2}$ with a step size of p1p-1, i.e., $u_{i}=\left\{\frac{i}{p-1}\mid i\in\{0,1,...,|u|\}\right\}$. This allows conversion to an integer problem (by multiplication of p1p-1) and apply Stars & Bars Theorem 2.
@@
Specifically, we set $k:=(n-1)^{2}$ and then define the set of sums that violate the constraint as $S:=\{s\in\mathbb{N}_{0}\mid 0\leq s\leq(n-2)(p-1)\}$. Thus |S|=(n2)(p1)|S|=(n-2)(p-1) We then compute the violations via:
@@
where $\binom{(n-1)^{2}}{m}$ accounts for the number of ways to choose $m$ out of $(n-1)^{2}$ variables that exceed p1p-1.
@@
$c2(n,p)=\sum_{s=0}^{(n-2)(p-1)}c_{2}(n,p,s)\\ =\sum_{s=0}^{|S|}\sum_{m=0}^{(n-1)^{2}}(-1)^{m}\binom{(n-1)^{2}}{m}\begin{cases}0&\hskip-40.00006pt\text{if }s-mp\leq 0,\\ \binom{s-mp+n^{2}-2n}{n^{2}-2n}&\text{else}.\end{cases}$ (18)
where |S|=(n2)(p1)|S|=(n-2)(p-1).